OpenStudy (anonymous):

{I'll give a medal and fan you :) } 4. Suppose y varies directly with x, and y- 12 when x= -3. What is the value of y when x=6 a. -12 b. 2 c. -2 d- 4

OpenStudy (anonymous):

@TheSmartOne could you help me pleasee?

OpenStudy (anonymous):

@shortycme could you help mee please?

OpenStudy (anonymous):

i think C

OpenStudy (anonymous):

thank you <3

OpenStudy (anonymous):

your welcome if you have anymore fill free to ask <3

OpenStudy (anonymous):

What is the equation of a line with a slope of -7 and a y-intercept of 6 ? a. 6y= -7x b. y=6x + (-7) c. -7y = x + 6 d. y= -7x + 6

OpenStudy (anonymous):

I do :)

OpenStudy (anonymous):

D

OpenStudy (anonymous):

is that all the questions you have

OpenStudy (anonymous):

noo. I don't

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

okay let me know if you do and i will try my best to help

OpenStudy (anonymous):

@shortycme

OpenStudy (anonymous):

either B or C

OpenStudy (anonymous):

D